OpenStudy (littlet):

How do you calculate the value of Cot 31 on a scientific calculator?

OpenStudy (jdoe0001):

\(\bf cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)}\qquad thus\implies cot(31^o)=\cfrac{cos(31^o)}{sin(31^o)}\)

OpenStudy (littlet):

Is that just Restating Cos is 31?

OpenStudy (littlet):

@jdoe0001

OpenStudy (littlet):

@phi can you help me?

OpenStudy (phi):

cotangent is the "flip" of tangent I would find the tangent of 31 : tan(31) then "invert it" using the \(x^{-1} \) key if you have one, or 1/tan(31) if you don't

OpenStudy (littlet):

so divide 1 by tan31?

OpenStudy (jdoe0001):

ohh hmmm w0t?

OpenStudy (littlet):

I got 2.3559

OpenStudy (phi):

is your calculator broken? what do you get for tan31 ?

OpenStudy (littlet):

oh for Tan 31 i got 0.600

OpenStudy (jdoe0001):

\(\bf cot(\theta)=\cfrac{1}{tan(\theta)}\impliedby \textit{make sure your calculator is in Degrees mode}\) notice, 31 is for degrees, not 31 radians, thus

OpenStudy (phi):

now what is 1 divided by that?

OpenStudy (littlet):

i now got 1.664

OpenStudy (phi):

yes, that looks better

OpenStudy (littlet):

alright lol. is that it?

OpenStudy (phi):

yes. also, up above joe showed another way to do this tan= sin/cos and if we flip both sides we get cot = cos/sin that means do cos(31) and then divide by sin(31) that gives you cotangent(31)

OpenStudy (littlet):

oh ok.

OpenStudy (phi):

you can try it (get practice using your calculator) and it should give the same number 1.664

OpenStudy (littlet):

@Phi could you help me with another one?

OpenStudy (phi):

please make a new post.

OpenStudy (littlet):

alright.